OpenStudy (anonymous):

Im so lost Can someone please help me with this geomtric mean.. ive started the equation but how to you find the square root? Question BELOW

OpenStudy (anonymous):

\[\frac{ 5 }{ x }\frac{ x }{ 12} x^2=60 x=\sqrt{60}\]

hartnn (hartnn):

huh ? do u just need square root of 60 ?

OpenStudy (anonymous):

yes how to you get the square root?

hartnn (hartnn):

can you factor 60 ?

OpenStudy (anonymous):

what does that mean

hartnn (hartnn):

like 10=2*5 24=2*2*2*3

hartnn (hartnn):

60=...?

OpenStudy (anonymous):

i dont ge it how did you get those numbers

hartnn (hartnn):

divide 60 by 2, what u get ?

OpenStudy (anonymous):

30

hartnn (hartnn):

so, 60 = 2*30 now 30 is a prime number ? no. so again divide 30 by 2, what u get ?

OpenStudy (anonymous):

15

hartnn (hartnn):

so, 60 = 2*2*15 now 15 is a prime number ? no. so again divide 15 by 2, can't divide.... divide 15 by 3, what u get ?

OpenStudy (anonymous):

5

hartnn (hartnn):

so, 60 = 2*2*3*5 right ? (since, 5 is a prime number, we stop) are you getting all this ?

OpenStudy (anonymous):

yes right !! & yes i am

hartnn (hartnn):

now comes the easy part ... \[\sqrt{x.x.y}=\sqrt{x.x}\sqrt y \] so what about \[\sqrt{60}=\sqrt{2.2.3.5}=...?\]

hartnn (hartnn):

i meant \[\sqrt{x.x.y}=\sqrt{x.x}\sqrt y=x\sqrt y\]

OpenStudy (anonymous):

wow sorry im horrible at math but is it \[2\sqrt{15}\]

hartnn (hartnn):

yes, its absolutely correct :) and you are not horrible at math, you understood all this, means you are good :)

hartnn (hartnn):

any doubts ?

OpenStudy (anonymous):

No Doubts!! & Thannk you so much for helping

hartnn (hartnn):

welcome ^_^